### v3

# Complete $k$-partite entanglement measure

Abstract
The $k$-partite entanglement, which focus on at most how many particles in the global system are entangled but separable from other particles, is complementary to the $k$-entanglement that reflects how many splitted subsystems are entangled under partitions of the systems in characterizing multipartite entanglement. Very recently, the theory of the complete $k$-entanglement measure has been established in [Phys. Rev. A 110, 012405 (2024)]. Here we investigate whether we can define the complete measure of the $k$-partite entanglement. Consequently, with the same spirit as that of the complete $k$-entanglement measure, we present the axiomatic postulates that a complete $k$-partite entanglement measure should require. Furthermore, we present two classes of $k$-partite entanglement measures and show that one is complete while the other one is unified but not complete except for the case of $k=2$.
Keywords: $k$-partite entanglement, $k$-partite entanglement measure, Unified/Complete $k$-partite entanglement measure
1 Introduction
In 2005, Gühne et al. introduced the $k$-partite entanglement in Ref. [1]. It is closely related to the $k$-producible state: if a quantum state is not $k$-producible, it is termed $(k+1)$-partite entangled. While the $k$-entanglement reflects how many splitted subsystems are entangled under partitions of the systems [2, 3], the $k$-partite entanglement concentrate on at most how many particles in the global system are entangled but separable from other particles [1, 4, 5]. It has been shown that $k$-producibility plays a crucial role in both quantum nonlocality [6, 7, 8, 9] and quantum metrology [10]. Particularly, $k$-producibly entangled states for larger $k$ exhibit higher sensitivity in phase estimation [11, 12, 13].
Clearly, these two ways of exhibiting entanglement, i.e., the $k$-entanglement and the $k$-partite entanglement, are complementary to each other in characterizing the multipartite entanglement which remains challenging to understand undeniably since the complexity increases substantially with the number of parties [14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Recently, the $k$-partite entanglement measure based on concurrence have been presnted [4, 5]. Very recently, we established the theory of the complete $k$-entanglement measure in Ref. [2]. It was shown that, in the framework of the complete measure of quantum correlation, the distribution of the correlation could be depicted exhaustively since the correlation could be compared not only between the global system and the subsystem (or the systems under arbitrary partition) but also between different subsystems (or the systems under arbitrary partition) [2, 21, 23, 24, 25, 26]. Along this line, the aim of this paper is to discuss how can we establish the axiomatic postulates for the complete $k$-partite entanglement measure.
The rest of the paper is arranged as follows. We review the concept of $k$-partite entanglement, the $k$-partite entanglement measures proposed in Ref. [4, 5], and the coarsening relation of multipartite partitions in Sec. 2. In Sec. 3, we present the definition of the complete $k$-partite entanglement measure, and then give two general ways of constructing $k$-partite entanglement measures and discuss whether they are complete in Sec. 4. Sec. 5 lists some examples of $k$-partite entanglement measures according to the methods in Sec. 4. Finally, in Sec. 6, we summarize the results of the paper.
2 Notations and Preliminaries
For convenience of discussing the complete measure of the $k$-partite entanglement in the next sections, we review some basic notations and terminologies in Sec 2.1, and introduce the $k$-partite entanglement measures proposed in literature so far in Sec. 2.2. We then introduce the coarsening relation of the multipartite partitions which is necessary when we discuss the completeness of a multipartite quantum correlation measure (also see in Ref. [23, 24]).
We fix some notations first. We denote by $A_{1}A_{2}\cdots A_{n}$ an $n$-partite quantum system. Let $X_{1}|X_{2}|\cdots|X_{m}$ be an $m$-partition of $A_{1}A_{2}\cdots A_{n}$ (for instance, partition $AB|C|DE$ is a $3$-partition of the 5-particle system $ABCDE$ with $X_{1}=AB$, $X_{2}=C$ and $X_{3}=DE$. The case of $m=n$ is just the original $n$-particle system without any other partition, namely, $A_{1}A_{2}\cdots A_{n}$ means $A_{1}|A_{2}|\cdots|A_{n}$. So $m<n$ in general unless otherwise specified). We denote by $\varDelta(X_{t})$ the number of subsystems contained in $X_{t}$, for instance, for the 3-partition $AB|C|DE$ of $ABCDE$, $\varDelta(X_{1})=\varDelta(AB)=2$, $\varDelta(X_{2})=\varDelta(C)=1$ and $\varDelta(X_{3})=\varDelta(DE)=2$. If $\varDelta(X_{t})\leqslant k$ for any $1\leqslant t\leqslant m$, we call it a $k$-fineness partition. We denote by $\Gamma_{\!k}^{f}$ the set of all $k$-fineness partitions of the given system $A_{1}A_{2}\cdots A_{n}$.
2.1 $k$-partite entanglement
A pure state $|\psi\rangle$ of an $n$-partite system $A_{1}A_{2}\cdots A_{n}$ with state space $\mathcal{H}^{A_{1}A_{2}\cdots A_{n}}$ is called $k$-producible ($1\leqslant k\leqslant n-1$), if it can be represented as [1]
$|\psi\rangle=|\psi\rangle^{X_{1}}|\psi\rangle^{X_{2}}\cdots|\psi\rangle^{X_{m}}$ (1)
under some $k$-fineness partition $X_{1}|X_{2}|\cdots|X_{m}$ of $A_{1}A_{2}\cdots A_{n}$. Let $\mathcal{S}^{X}$ be the set of all density operators acting on the state space $\mathcal{H}^{X}$. For mixed state $\rho\in\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$, if it can be written as a convex combination of $k$-producible pure states, i.e., $\rho=\sum\limits_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$ with $|\psi_{i}\rangles$ are $k$-producible, it is called $k$-producible, where the pure state $|\psi_{i}\rangles$ might be $k$-producible in different $k$-fineness partitions. If a quantum state is not $k$-producible, it is termed $(k+1)$-partite entangled. Note that if $|\psi\rangle$ admits the form as in Eq. (1), it is called $m$-separable. $|\psi\rangle$ is $m$-entangled if it is not $m$-separable. An $n$-partite mixed state $\rho$ is $m$-separable if it can be written as $\rho=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$ with $|\psi_{i}\rangles$ are $m$-separable, wherein the contained $\{|\psi_{i}\rangle\}$ can be $m$-separable with respect to different $m$-partitions. Otherwise, it is called $m$-entangled. By definition, the $k$-partite entanglement is different from the $k$-entanglement in general, but they are equivalent only in some special cases. For example, the $n$-partite entangled state is just the genuine multipartite entangled state and the one-producible state coincides with the fully separable state. If $|\psi\rangle^{ABC}$ is a genuine entangled state, then $|\psi\rangle^{ABC}|\psi\rangle^{D}|\psi\rangle^{E}|\psi\rangle^{F}$ is four-separable and three-partite entangled state. Also note that, a state of which some reduced state of $m$ parties is genuinely entangled, contains $m$-partite entanglement, but not vice versa in general [1]. For more clarity, we compare $3$-partite entangled pure state with $3$-entangled pure state in Fig. 1.
Figure 1: (color online). (a) 33-partite entangled pure state $|\Psi\rangle=|\psi\rangle^{A^{\prime}_{1}A^{\prime}_{2}A^{\prime}_{3}}|\psi\rangle^{B^{\prime}_{1}B^{\prime}_{2}\cdots B^{\prime}_{p}}\cdots|\psi\rangle^{X^{\prime}_{1}X^{\prime}_{2}\cdots X^{\prime}_{q}}|\phi\rangle^{A_{1}A_{2}}|\phi\rangle^{B_{1}B_{2}}\cdots|\phi\rangle^{X_{1}X_{2}}|\varphi\rangle^{A}|\varphi\rangle^{B}\cdots$ $|\varphi\rangle^{X}$, where $|\psi\rangle^{A^{\prime}_{1}A^{\prime}_{2}A^{\prime}_{3}}$, $|\psi\rangle^{B^{\prime}_{1}B^{\prime}_{2}\cdots B^{\prime}_{p}}$, $\dots$, $|\psi\rangle^{X^{\prime}_{1}X^{\prime}_{2}\cdots X^{\prime}_{q}}$ are genuinely entangled states, $3\leqslant p\leqslant q$, $|\phi\rangle^{A_{1}A_{2}}$, $|\phi\rangle^{B_{1}B_{2}}$, $\dots$, $|\phi\rangle^{X_{1}X_{2}}$ are entangled states. In fact, if one of $|\psi\rangle^{A^{\prime}_{1}A^{\prime}_{2}A^{\prime}_{3}}$, $|\psi\rangle^{B^{\prime}_{1}B^{\prime}_{2}\cdots B^{\prime}_{p}}$, $\dots$, $|\psi\rangle^{X^{\prime}_{1}X^{\prime}_{2}\cdots X^{\prime}_{q}}$ is genuinely entangled, $|\Psi\rangle$ is also 33-partite entangled. Here we just take the general form of a $3$-partite entangled pure state. (b) $|\Phi\rangle=|\psi\rangle^{A_{1}A_{2}\cdots A_{k}}|\psi\rangle^{B_{1}B_{2}\cdots B_{l}}$ with $k$, $l\geqslant 0$, $k+l\geqslant 3$, is a 33-entangled pure state if one of the following is true: (i) $|\psi\rangle^{A_{1}A_{2}\cdots A_{k}}$ and $|\psi\rangle^{B_{1}B_{2}\cdots B_{l}}$ are genuinely entangled states, $k$, $l\geqslant 3$, (ii) $|\psi\rangle^{A_{1}A_{2}\cdots A_{k}}$ and $|\psi\rangle^{B_{1}B_{2}\cdots B_{l}}$ are entangled states, $k=l=2$, (iii) If $k=0$ or $l=0$, $|\Phi\rangle$ is genuinely entangled.
A pure state $|\phi\rangle$ is said to be genuinely $k$-producible (or genuinely $k$-partite entangled) [1] if it is $k$-producible but not (k1)(k-\!1)-producible. A mixed state $\rho\in\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$ is genuinely $k$-producible if it is $k$-producible and for any $k$-producible pure states ensemble of $\rho$, $\rho=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$, there is at least one $|\psi_{i}\rangle$ is genuinely $k$-producible.
Let $\mathcal{S}_{P(k)}$ $(k=1,2,\ldots,n-1)$ denote the set of $k$-producible quantum states in $\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$ and $\mathcal{S}_{P(n)}:=\mathcal{S}$. It follows that
$\mathcal{S}_{P(1)}\subset\mathcal{S}_{P(2)}\subset\cdots\subset\mathcal{S}_{P(n-1)}\subset\mathcal{S}_{P(n)},$ (2)
$\mathcal{S}\setminus\!\mathcal{S}_{P(k)}$ is the set consisting of all $(k+1)$-partite entangled states, and $\mathcal{S}_{P(k)}\!\!\setminus\!\mathcal{S}_{P(k-1)}$ is the set of all genuinely $k$-producible states.
2.2 $k$-partite entanglement measures via $q$-concurrence and $\alpha$-concurrence
A positive function $E_{(k)}:\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}\longrightarrow\mathbb{R}_{+}$ is called a $k$-partite entanglement measure ($k$-PEM) if it fulfills: (i) $E_{(k)}(\rho)=0$ for any $\rho\in\mathcal{S}_{P(k-1)}$ and $E_{(k)}(\rho)>0$ for any $\rho\in\mathcal{S}\setminus\mathcal{S}_{P(k-1)}$, (ii) $E_{(k)}(\rho)$ does not increase under $n$-partite local operations and classical communication (LOCC), namely, $E_{(k)}(\varepsilon(\rho))\leq E_{(k)}(\rho)$ for any $n$-partite LOCC $\varepsilon$. Item (ii) guarantees that $E_{(k)}$ is invariant under local unitary operation. In addition, a $k$-PEM $E_{(k)}$ on $\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$ is convex and non-increasing on average under $n$-partite LOCC, it is call a $k$-partite entanglement monotone ($k$-PEMo). A $k$-PEM/$k$-PEMo $E_{(k)}$ is called a genuine $k$-PEM/$k$-PEMo if $E_{(k)}(\rho)>0$ but $E_{(k+1)}(\rho)=0$ for any genuinely $k$-partite entangled state $\rho$.
Hong et al. presented a $k$-PEMo in Ref. [4] via concurrence. For any pure state $|\psi\rangle\in\mathcal{H}^{A_{1}A_{2}\cdots A_{n}}$, the $k$-PEMo was defined as [4]
$C_{(k)}(|\psi\rangle)=\min\limits_{\Gamma_{\!k-\!1}^{f}}\frac{\sum_{t=1}^{m}\sqrt{2[1-{\rm Tr}(\rho_{X_{t}}^{2})]}}{m},$ (3)
where $\rho_{X_{t}}={\rm Tr}_{{\overline{X}}_{t}}(|\psi\rangle\langle\psi|)$, ${\overline{X}}_{t}$ is the complement of subsystem $X_{t}$, the minimum is taken over all the (k1)(k\!-\!1)-fineness partitions in $\Gamma_{\!k-\!1}^{f}$. For mixed states, it is defined by the convex-roof extension, i.e.,
$C_{(k)}(\rho)=\min\limits_{p_{i},|\psi_{i}\rangle}\sum_{i}p_{i}C_{(k)}(|\psi_{i}\rangle),$
where the minimum runs over all ensembles $\{p_{i},|\psi_{i}\rangle|\rho=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|\}$. In what follows, we give only the measures for pure states, for the case of mixed states they are all defined by the convex-roof extension with no further statement. Obviously, any measure that is defined in this way is convex.
Very recently, Li et al. proposed two $k$-PEMos in Ref. [5]. For any pure state $|\psi\rangle\in\mathcal{H}^{A_{1}A_{2}\cdots A_{n}}$, the $k$-PEMo via the $q$-concurrence was defined as [5]
$C_{q(k)}(|\psi\rangle)=\min\limits_{\Gamma_{\!k-\!1}^{f}}\sqrt{\frac{\sum_{t=1}^{m}[1-{\rm Tr}(\rho_{X_{t}}^{q})]}{m}},$ (4)
and the $k$-PEMo via the $\alpha$-concurrence was expressed by [5]
$C_{\alpha(k)}(|\psi\rangle)=\min\limits_{\Gamma_{\!k-\!1}^{f}}\sqrt{\frac{\sum_{t=1}^{m}[{\rm Tr}(\rho_{X_{t}}^{\alpha})-1]}{m}},$ (5)
where the minimum is taken over all the (k1)(k\!-\!1)-fineness partitions in $\Gamma_{\!k-\!1}^{f}$. Note here that, the notations here are different from $E_{q-k}$ and $E_{\alpha-k}$ in Ref. [5] ($C_{q(k+1)}=E_{q-k}$, $C_{\alpha(k+1)}=E_{\alpha-k}$). Ref. [5] also gave the following two $k$-PEMos:
$C_{G,q(k)}(|\phi\rangle)=\left(\frac{\prod\limits_{\gamma_{i}\in\Gamma_{\!k-\!1}^{f}}\left[\sum\limits_{t=1}^{m_{i}}(1-{\rm Tr}\rho_{X_{t(i)}}^{q})\right]}{\prod_{i=1}^{\left|\Gamma_{\!k-\!1}^{f}\right|}m_{i}}\right)^{\frac{1}{2\left|\Gamma_{\!k-\!1}^{f}\right|}}$ (6)
and
$C_{G,\alpha(k)}(|\phi\rangle)=\left(\frac{\prod\limits_{\gamma_{i}\in\Gamma_{\!k-\!1}^{f}}\left[\sum\limits_{t=1}^{m_{i}}({\rm Tr}\rho_{X_{t(i)}}^{\alpha}-1)\right]}{\prod_{i=1}^{\left|\Gamma_{\!k-\!1}^{f}\right|}m_{i}}\right)^{\frac{1}{2\left|\Gamma_{\!k-\!1}^{f}\right|}},$~~~ (7)
where $\rho_{X_{t(i)}}$ is the reduced density operator with respect to subsystem $X_{t(i)}$}, and $m_{i}$ refers to $\gamma_{i}$ is a $m_{i}$-partition, $\left|\Gamma_{\!k-\!1}^{f}\right|$ is the cardinal number of $\Gamma_{\!k-\!1}^{f}$. The notations in Eqs. (6), (7) are different from $\varepsilon_{q-k}$ and $\varepsilon_{\alpha-k}$ in Ref. [5] ($C_{G,q(k+1)}=\varepsilon_{q-k}$, $C_{G,\alpha(k+1)}=\varepsilon_{\alpha-k}$).
2.3 Coarsening relation of multipartite partitions
Let $X_{1}|X_{2}|\cdots|X_{k}$ and $Y_{1}|Y_{2}|\cdots|Y_{l}$ be two partitions of $A_{1}A_{2}\cdots A_{n}$ or subsystem of $A_{1}A_{2}\cdots A_{n}$, $k\leq n$, $l\leq n$. We denote by [2, 23]
$X_{1}|X_{2}|\cdots|X_{k}\succ^{a}Y_{1}|Y_{2}|\cdots|Y_{l},$ (8)
$X_{1}|X_{2}|\cdots|X_{k}\succ^{b}Y_{1}|Y_{2}|\cdots|Y_{l},$ (9)
$X_{1}|X_{2}|\cdots|X_{k}\succ^{c}Y_{1}|Y_{2}|\cdots|Y_{l}$~ (10)
if $Y_{1}|Y_{2}|\cdots|Y_{l}$ can be obtained from $X_{1}|X_{2}|\cdots|X_{k}$ by
  • (a)
    Discarding some subsystem(s) of $X_{1}|X_{2}|\cdots|X_{k}$,
  • (b)
    Combining some subsystems of $X_{1}|X_{2}|\cdots|X_{k}$,
  • (c)
    Discarding some subsystem(s) of some subsystem(s) $X_{t}$ provided that $X_{t}=A_{t(1)}A_{t(2)}\cdots A_{t(f(t))}$ with $f(t)\geqslant 2$, $1\leqslant t\leqslant k$,
respectively. For example,
$A|B|C|D\succ^{a}A|B|D\succ^{a}B|D,$
$A|B|C|D\succ^{b}AC|B|D\succ^{b}AC|BD,$
$A|BC\succ^{c}A|B.$
Here, we denote $ABCD$, ABDABD, BDBD and $AB$ by $A|B|C|D$, $A|B|D$, $B|D$ and $A|B$, respectively.
For any subsystem of $A_{1}A_{2}\cdots A_{n}$ with arbitrary partition, it can always be derived from the global system via the coarsening relations (a)-(c) or some of them. So, based on these three coarsening relations, we can analyze not only the information relation between any subsystem and the global subsystem but also the information relation between any subsystems in detail. For instance, based on these coarsening relations, we have established the complete global entanglement measure [21, 24], the complete genuine entanglement measure [23, 24], the complete multipartite quantum discord [25], the complete multipartite quantum mutual information [26] and the complete $k$-entanglement measure [2]. Furthermore, we discussed the complete monogamy relation of these measures [2, 21, 23, 24, 25, 26], where exploring the monogamy relation of the quantum correlations is one of the fundamental tasks in the quantum resource theory [21, 25, 27, 28, 29, 30, 31, 32, 33, 34].
3 Completeness of the $k$-PEM
When we deal with the various quantum correlations living in a multipartite system, the most quintessential relation in a multipartite system is indeed the the coarsening relation, i.e., the coarsening relations of type (a)-(c). The “completeness” of a measure for multipartite quantum correlation mainly refers to that there is a unified criterion for quantifying different subsystems or systems under different partition, which means the amount of the quantum correlations contained in different particles or particles under arbitrary partition can be compared with each other consistently and compatibly [2, 21, 23, 25, 26]. It makes up for the previous bipartite measure which can only quantify the quantum correlation under the given bipartite splitting. By reviewing the key point in defining a complete measure of quantum correlation [2, 21, 23, 25, 26], we can conclude that there are two steps to reveal such a completeness of a given measure: the first step is the unification condition which is mainly related to the coarsening relation of type (a), and the second one is the hierarchy condition which is determined by the coarsening relation of type (b).
We now give the axiomatic postulates for the unified $k$-PEM and the complete $k$-PEM based on the coarsening relation of the partitions of the system. Hereafter, $E_{(k)}(X)$ denotes $E_{(k)}(\rho^{X})$. A $k$-PEM $E_{(k)}$ is called unified if it satisfies the unification condition: (i) (symmetry) $E_{(k)}({A_{1}A_{2}\cdots A_{n}})=E_{(k)}({A_{\pi(1)}A_{\pi(2)}\cdots A_{\pi(n)}})$ for all $\rho\in\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$ and any permutation $\pi$ of $\{1,2,\cdots,n\}$; (ii) (additivity) $E_{(k)}(A_{1}A_{2}\cdots A_{r}\otimes A_{r+1}A_{r+2}\cdots A_{n})=E_{(k)}(A_{1}A_{2}\cdots A_{r})+E_{(k)}(A_{r+1}A_{r+2}\cdots A_{n})$ holds for all $\rho^{A_{1}A_{2}\cdots A_{r}}\otimes\rho^{A_{r+1}A_{r+2}\cdots A_{n}}$; (iii) ($k$-monotone)
$E_{(k)}(A_{1}A_{2}\cdots A_{n})\leqslant E_{(k-1)}(A_{1}A_{2}\cdots A_{n})$ (11)
holds for all $\rho\in\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$, $k\geqslant 3$; and (iv) (coarsening monotone)
$E_{(k)}(X_{1}|X_{2}|\cdots|X_{p})\geqslant E_{(k)}(Y_{1}|Y_{2}|\cdots|Y_{q})$ (12)
[…]
A unified $k$-PEM $E_{(k)}$ is called complete if it satisfies the hierarchy condition additionally: (v) (tight coarsening monotone)
$E_{(k)}(A_{1}A_{2}\cdots A_{n})\geqslant E_{(k)}(Y_{1}|Y_{2}|\cdots|Y_{q})$ (13)
holds for all state $\rho\in\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$ whenever $A_{1}A_{2}\cdots A_{n}\succ^{b}Y_{1}|Y_{2}|\cdots|Y_{q}$ such that, for any $i$, either $\rho^{Y_{i}}$ is pure or $\rho^{Y_{i}}$ is the reduced state of some genuinely entangled pure state (or entangled bipartite pure state), $1\leqslant i\leqslant q<n$. If a $k$-PEM $E_{(k)}$ satisfies Eq. (13), we call it is tightly coarsening monotonic. For example, take $|\psi\rangle=|\psi\rangle^{ABCD}|\psi\rangle^{EF}|\psi\rangle^{GH}|\psi\rangle^{I}$ with $|\psi\rangle^{ABCD}$ is genuinely entangled, $|\psi\rangle^{EF}$ and $|\psi\rangle^{GH}$ are entangled, then the partition $AB|CD|EF|GHI$ is such a case since $|\psi\rangle^{EF}$ and $|\psi\rangle^{GHI}$ are pure states, $\rho^{AB}$ and $\rho^{CD}$ are reduced states of the genuine pure state $|\psi\rangle^{ABCD}$. That is, at least intuitively, such a coarsening operation can not generate $k$-partite entanglement. So we hope Eq. (13) should be satisfied. In addition, we take
$|\psi\rangle=|\psi\rangle^{ABC}|\psi\rangle^{DE}|\psi\rangle^{F}|\psi\rangle^{GH}|\psi\rangle^{IJ},$
then
$E_{(3)}(|\psi\rangle)=E_{(3)}(|\psi\rangle^{ABC}).$
But
$E_{(3)}(A|B|C|D|EFGI|H|J)=E_{(3)}(ABC)+E_{(3)}(D|EFGI|H|J)$
which is larger than $E_{(3)}(|\psi\rangle)$ whenever $|\psi\rangle^{DE}$, $|\psi\rangle^{GH}$ and $|\psi\rangle^{IJ}$ are entangled states. Namely, Eq. (13) is violated under the general coarsening operation of type (b). We thus adjust the hierarchy condition (i.e., the tight coarsening monotone condition) as Eq. (13) in stead of the “strict” hierarchy condition in Refs. [2, 21].
One need note here that, for any given $k$-PEM $E_{(k)}$,
$E_{(k)}(X_{1}|X_{2}|\cdots|X_{p})\geqslant E_{(k)}(X^{\prime}_{1}|X^{\prime}_{2}|\cdots|X^{\prime}_{p})$ (14)
holds for any $\rho\in\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$ whenever $X_{1}|X_{2}|\cdots|X_{p}\succ^{c}X^{\prime}_{1}|X^{\prime}_{2}|\cdots|X^{\prime}_{p}$ since $\rho^{X^{\prime}_{1}|X^{\prime}_{2}|\cdots|X^{\prime}_{p}}$ is obtained from $\rho^{X_{1}|X_{2}|\cdots|X_{p}}$ by a partial trace and such a partial trace is indeed a $p$-partite LOCC, $1\leqslant k\leqslant p<n$.
We take the 4-partite system $ABCD$ for example. $E_{(4)}$ is $k$-monotonic means
$E_{(4)}(ABCD)\leqslant E_{(3)}(ABCD)\leqslant E_{(2)}(ABCD)$
for any $\rho\in\mathcal{S}^{ABCD}$, and $E_{(3)}$ is coarsening monotonic refers to
$E_{(3)}(ABCD)\geqslant E_{(3)}(ABC),$
$E_{(3)}(ABCD)\geqslant E_{(3)}(ABD),$
$E_{(3)}(ABCD)\geqslant E_{(3)}(ACD),$
$E_{(3)}(ABCD)\geqslant E_{(3)}(BCD),$
for any state $\rho\in\mathcal{S}^{ABCD}$. If $E_{(3)}$ is tightly coarsening monotonic,
$E_{(3)}(ABCD)\geqslant E_{(3)}(A|B|CD),$
$E_{(3)}(ABCD)\geqslant E_{(3)}(A|BC|D),$
$E_{(3)}(ABCD)\geqslant E_{(3)}(AB|C|D),$
$E_{(3)}(ABCD)\geqslant E_{(3)}(AC|B|D),$
$E_{(3)}(ABCD)\geqslant E_{(3)}(AD|B|C),$
$E_{(3)}(ABCD)\geqslant E_{(3)}(A|C|BD)$
for any $\rho\in\mathcal{S}^{ABCD}$.
4 Two classes of $k$-PEMos
[…]
4.1 $k$-PEMo from unified multipartite entanglement measure
[…]
In general, for any given pure state $|\psi\rangle=|\psi\rangle^{A_{1}A_{2}\cdots A_{n}}$ in $\mathcal{H}^{A_{1}A_{2}\cdots A_{n}}$, we assume it is not (k1)(k-1)-producible. Then there exists a $l$-fitness partition $X_{1}|X_{2}|\cdots|X_{m}$, $l\geqslant k$, such that
$\left\{\begin{array}[]{l}\!\!\Delta(X_{t}):=s(t)\geqslant k,\\ \!\!\rho^{X_{t}}~\textrm{is a genuinely entangled pure state}\end{array}\right$.~~~~ (17)
for some subsystem $X_{t}$ in the partition $X_{1}|X_{2}|\cdots|X_{m}$. Let $t_{1}$, $t_{2}$, $\dots$, $t_{r}$ be all of the subscripts such that $X_{t_{i}}$ satisfies the condition (17) corresponding to all possible $l$-fitness partitions with $l\geqslant k$. It turns out that
$|\psi\rangle=|\psi\rangle^{X_{t_{1}}}|\psi\rangle^{X_{t_{2}}}\cdots|\psi\rangle^{X_{t_{r}}}|\phi\rangle^{X_{*}}$ (18)
under some permutation of the subsystems, where $X_{*}$ denotes the subsystem complementary to $X_{t_{1}}X_{t_{2}}\cdots X_{t_{r}}$. In such a sense, we can quantify the $k$-partite entanglement of $|\psi\rangle$ by
$E_{(k)}(|\psi\rangle)=\sum\limits_{j=1}^{r}E^{(s(t_{j}))}(|\psi\rangle^{X_{t_{j}}})$ (19)
for any given unified MEM $E^{(n)}$.
With the notations above, we give the following definition of a $k$-PEMo:
$E_{(k)}(|\psi\rangle)=\begin{cases}\sum\limits_{j=1}^{r}E^{(s(t_{j}))}(|\psi\rangle^{X_{t_{j}}}),&s(t_{j})\geq k~\text{for some}~j,\\ 0,&\text{otherwise}.\end{cases}$ (20)
Theorem 1.
Let $h$ be a reduced function. If $E^{(n)}(|\psi\rangle)=\frac{1}{2}\sum_{i}h(\rho^{A_{i}})$, then $E_{(k)}$ is a unified $k$-PEMo, and moreover, $E_{(k)}$ is a complete $k$-PEMo whenever $h$ is subadditve.
Proof.
Items (i) and (ii) are straightforward. For any $|\psi\rangle\in\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$, we suppose that $E_{(k)}(|\psi\rangle)=\sum\limits_{j=1}^{r}E^{(s(t_{j}))}(|\psi\rangle^{X_{t_{j}}})>0$ for some partition $X_{t_{1}}|X_{t_{2}}|\cdots|X_{t_{r}}|X_{*}$. It turns out that $E_{(k)}(|\psi\rangle^{X_{t_{j}}})=E_{(k-1)}(|\psi\rangle^{X_{t_{j}}})$ for any $j$, $E_{(k)}(|\phi\rangle^{X_{*}})=0$, but it is possible that $E_{(k-1)}(|\phi\rangle^{X_{*}})>0$, which implies (iii) is true. For any partition $X_{1}|X_{2}|\cdots|X_{p}$ of $A_{1}A_{2}\cdots A_{n}$, $p\geqslant k$, we consider $X_{2}|\cdots|X_{p}$ w.n.l.g., namely the partition that by discarding $X_{1}$ from $X_{1}|X_{2}|\cdots|X_{p}$. If $X_{1}$ is $X_{t_{j}}$ or some subsytem(ssome) of $X_{t_{j}}$, (iv) is clear since $E^{(s)}$ is unified [24] [a unified MEM is decreasing under the coarsening relation of type (a)]. If $X_{1}$ is $X_{*}$ or some subsytem(s) of $X_{*}$, (iv) is clear since $E_{(k)}(|\psi\rangle)=E_{(k)}(\rho^{X_{2}|\cdots|X_{p}})$. If $X_{1}$ cotains some subsytem(s) of some $X_{t_{j}}$, (iv) is also ture since $E^{(s)}$ is unified. The other cases can be argued similarly.
If the reduced function $h$ is subadditive, then $E^{(n)}$ is complete [21, 24], which means that $E^{(n)}$ is nonincreasing under the coarsening operation of type (b). This completes the proof. ∎
Another candidate for the unified global MEM is the one defined by the sum of all bipartite entanglement [22], i.e.,
$\qquad\mathcal{E}^{(n)}(|\psi\rangle^{A_{1}A_{2}\cdots A_{n}})=\left\{\begin{array}[]{ll}\frac{1}{2}\sum\limits_{i_{1}\leq\cdots\leq i_{s},s<n/2}h(\rho^{A_{i_{1}}A_{i_{2}}\cdots A_{i_{s}}}),&\mbox{if}~n~\mbox{is odd},\\ \frac{1}{2}\sum\limits_{i_{1}\leq\cdots\leq i_{s}<n,s\leq n/2}h(\rho^{A_{i_{1}}A_{i_{2}}\cdots A_{i_{s}}}),&\mbox{if}~n~\mbox{is even},\end{array}\right$.~~~~~ (23)
where $h$ is some given reduced function. We denote by $\mathcal{E}_{(k)}$ the quantity that is defined as in Eq. (20) just with $\mathcal{E}^{(s(t_{j}))}$ replacing $E^{(s(t_{j}))}$. Using similar arguments as in the proof of Theorem 1, we can conclude the following theorem.
Theorem 2.
$\mathcal{E}_{(k)}$ is a unified $k$-PEMo, and $\mathcal{E}_{(k)}$ is complete if $h$ is subadditive.
In Eq. (20), $|\psi\rangle^{X_{t_{j}}}$ is genuinely entangled, so $E^{(s(t_{j}))}$ can be any genuine entanglement measure instead. For example, the genuine entanglement measure from the minimal reduced function, which is defined by [24]
$\varepsilon_{g^{\prime\prime}}^{(n)}(|\psi\rangle^{A_{1}A_{2}\cdots A_{n}})=h(\rho^{X})$
where $h$ is some given reduced function, $X\subsetneq\{A_{1},A_{2},\dots,A_{n}\}$, namely the minimum runs over all possible reduced states. Then the corresponding $k$-PEMo is not unified in general since $\varepsilon_{g^{\prime\prime}}^{(n)}$ may increase under the coarsening relation of type (a) [24].
4.2 $k$-PEMo from the minimal sum
Let $|\psi\rangle=|\psi\rangle^{A_{1}A_{2}\cdots A_{n}}$ be a pure state in $\mathcal{H}^{A_{1}A_{2}\cdots A_{n}}$ and $h$ be a reduced function. For any $\gamma_{i}^{f}\in\Gamma_{k}^{f}$, we write
$\mathcal{P}_{k}^{\gamma_{i}^{f}}(|\psi\rangle)\equiv\frac{1}{2}\sum_{t=1}^{m}h(\rho^{X_{t(i)}}),~1\leqslant k<n,$ (24)
[…]
$E^{\prime}_{(k)}(|\psi\rangle)=\min\limits_{\Gamma_{\!k-\!1}^{f}}\mathcal{P}_{\!k-\!1}^{\gamma_{i}^{f}}(|\psi\rangle),$ (25)
where the minimum is taken over all feasible (k1)(k\!\!-\!\!1)-fineness partitions in $\Gamma_{\!k-\!1}^{f}$. By definition, for any $\rho\in\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$, $E^{\prime}_{(k)}(\rho)>0$ if and only if $\rho$ is $k$-partite entangled.
Theorem 3.
$E^{\prime}_{(k)}$ is a unified $k$-PEMo and $E^{\prime}_{(2)}$ is complete if the reduced function is subadditive.
Proof.
By definition, it is straightforward that $E^{\prime}_{(k)}$ is a $k$-PEMo. We show below it satisfies the unification conditions (i)-(iv). We only need to check items (iii) and (iv) since (i) and (ii) are clear. Since $\Gamma^{f}_{k\!-\!2}\subseteq\Gamma^{f}_{k\!-\!1}$, this implies (iii) is true. For any given $|\psi\rangle\in\mathcal{H}^{A_{1}A_{2}\cdots A_{n}}$, we assume w.n.l.g. that
$E^{\prime}_{(k)}(|\psi\rangle)=\frac{1}{2}\left[h(\rho^{X_{1}})+h(\rho^{X_{2}})+h(\rho^{X_{3}})\right]$
for some $X_{1}|X_{2}|\cdots|X_{m}$. If
$Y_{1}|Y_{2}|\cdots|Y_{p}\succ^{a}Y^{\prime}_{1}|Y^{\prime}_{2}|\cdots|Y^{\prime}_{q}$
with $Y^{\prime}_{1}|Y^{\prime}_{2}|\cdots|Y^{\prime}_{q}$ is obtained by discarding subsystem(s) $Y_{i}$s such that $Y_{i}$s are contained in $X_{4}X_{5}\cdots X_{m}$, Eq. (12) is clear. If $Y^{\prime}_{1}|Y^{\prime}_{2}|\cdots|Y^{\prime}_{q}$ is obtained by discarding subsystem(s) $Y_{i}$s such that $Y_{i}$s are contained in $X_{1}X_{2}X_{3}$, there are three subcases: (a) $Y_{i}=X_{1}$ (w.n.l.g.), (b) $Y_{i}=X_{21}$, $X_{2}=X_{21}X_{22}$ (w.n.l.g.), and (c) $Y_{i}=X_{1}X_{21}$ (w.n.l.g.). For the subcase of (a), it turns out that
$E^{\prime}_{(k)}(|\psi\rangle^{X_{1}X_{2}X_{3}})$
$=$ $\frac{1}{2}\left[h(\rho^{X_{1}})+h(\rho^{X_{2}})+h(\rho^{X_{3}})\right]$
$>$ $\frac{1}{2}\sum_{i}p_{i}\left[h(\rho^{X_{2(i)}})+h(\rho^{X_{3(i)}})\right]$
$\geqslant$ $E^{\prime}_{(k)}(\rho^{X_{2}X_{3}})$
for any pure state ensemble $\{p_{i},|\psi\rangle^{X_{2(i)}X_{3(i)}}\}$ of $\rho^{X_{2}X_{3}}$ since $h$ is concave. That is, Eq. (12) holds true still. For the subcase of (b), it follows that
$\quad E^{\prime}_{(k)}(|\psi\rangle^{X_{1}X_{2}X_{3}})\geqslant\frac{1}{2}\sum_{i}p_{i}\left[h(\rho^{X_{1(i)}})+h(\rho^{X_{22(i)}})+h(\rho^{X_{3(i)}})\right]\geqslant E^{\prime}_{(k)}(\rho^{X_{12}X_{2}X_{3}})$
for any pure state ensemble $\{p_{i},|\psi\rangle^{X_{1(i)}X_{22(i)}X_{3(i)}}\}$ of $\rho^{X_{1}X_{22}X_{3}}$, where the first inequality holds since discarding $X_{11}$ is a partial trace which is a special LOCC on $X_{1}|X_{2}|X_{3}$, and $E^{\prime}_{(k)}$ in such a case is $E^{(3)}$ acting on $\mathcal{S}^{X_{1}X_{2}X_{3}}$ which leads to decreasing under LOCC. The subcase of (c) is clear from the arguments for (a) together with (b). That is, Eq. (12) is valid for both subcases.
It is clear that the completeness of $E^{\prime}_{(2)}$ is reduced to the subadditivity of the reduced function. This completes the proof. ∎
By definition, if the reduced function is subadditive, it can be easily checked that
$E^{\prime}_{(k)}(\rho)\leqslant E_{(k)}(\rho)\leqslant\mathcal{E}_{(k)}(\rho).$ (26)
If $h$ is subadditive, then the minimal partition is the ones that contained in $\Gamma_{\!k-\!1}^{f}\!\!\setminus\!\Gamma_{\!k-\!2}^{f}$. $E^{\prime}_{(k)}$ is not complete in general if $k\geqslant 3$. For example, if
$\qquad\quad E^{\prime}_{(3)}(|\psi\rangle^{ABCD}|\psi\rangle^{EF}|\psi\rangle^{GHI})=\frac{1}{2}\left[h(\rho^{AB})+h(\rho^{CD})+h(\rho^{GH})+h(\rho^{I})\right],$
it follows that
$E^{\prime}_{(3)}(A|BC|D|E|F|G|H|I)$
$=$ $\frac{1}{2}\left[h(\rho^{A})+h(\rho^{BCD})+h(\rho^{GH})+h(\rho^{I})\right]$
$\geqslant$ $E^{\prime}_{(3)}(ABCDEFGH)$
whenever $h(\rho^{D})\geq h(\rho^{A})>h(\rho^{AB})$.
It can be easily checked that $C_{(k)}$ in Eq. (3), $C_{q(k)}$, $C_{\alpha(k)}$, $C_{G,q(k)}$ and $C_{G,\alpha(k)}$ in Eqs. (4)-(7) are not unified. Let
$\qquad\quad C_{(3)}(|\psi\rangle^{ABC}|\psi\rangle^{DE}|\psi\rangle^{FGH})=\frac{1}{5}\left[h(\rho^{AB})+h(\rho^{C})+h(\rho^{FG})+h(\rho^{H})\right],$
then $C_{(2)}(|\psi\rangle^{ABC}|\psi\rangle^{DE}|\psi\rangle^{FGH})=\frac{1}{8}(h_{A}+h_{B}+h_{C}+h_{D}+h_{E}+h_{F}+h_{G}+h_{H})$. Hereafter, we denote $h(\rho^{X})$ by $h_{X}$ for simplicity. Clearly, it is not necessary that $C_{(3)}(|\psi\rangle^{ABC}|\psi\rangle^{DE}|\psi\rangle^{FGH})\leq C_{(2)}(|\psi\rangle^{ABC}|\psi\rangle^{DE}|\psi\rangle^{FGH})$. So it is not $k$-monotonic. Let
$\qquad\quad C_{(3)}(|\psi\rangle^{AB}|\psi\rangle^{C}|\psi\rangle^{DEF})=\frac{1}{4}(h_{AB}+h_{C}+h_{DE}+h_{F})=\frac{1}{4}(h_{DE}+h_{F}).$
Then $C_{(3)}(|\psi\rangle^{AB}|\psi\rangle^{DEF})=\frac{1}{3}(h_{AB}+h_{DE}+h_{F})=\frac{1}{3}(h_{DE}+h_{F})$, i.e., $C_{(3)}$ is not coarsening monotonic. In addition $C_{(3)}(|\psi\rangle^{AB}|\psi\rangle^{C}+C_{(3)}(|\psi\rangle^{DEF})=C_{(3)}(|\psi\rangle^{DEF})=\frac{1}{2}(h_{DE}+h_{F})$, so $C_{(3)}$ is not additive. For $|\psi\rangle^{AB}|\psi\rangle^{CD}$, $C_{(2)}(|\psi\rangle^{AB}|\psi\rangle^{CD})=\frac{1}{4}(h_{A}+h_{B}+h_{C}+h_{D})=\frac{1}{2}(h_{A}+h_{C})$ may be not larger than $C_{(2)}(AB|C|D)=2h_{C}/3$.
Take $|\psi\rangle^{AB}|\psi\rangle^{C}|\psi\rangle^{DE}$, then
$C_{q(2)}(|\psi\rangle^{AB}|\psi\rangle^{C}|\psi\rangle^{DE})=\sqrt{\frac{2}{5}}\sqrt{h_{A}+h_{D}}$
$\neq$ $C_{q(2)}(|\psi\rangle^{AB})+C_{q(2)}(|\psi\rangle^{C}|\psi\rangle^{DE})=\sqrt{h_{A}}+\sqrt{2h_{D}/3},$
in general. So it is not additive. For $|\psi\rangle^{AB}|\psi\rangle^{C}|\psi\rangle^{DEF}$, we assume that
$C_{q(3)}(|\psi\rangle^{AB}|\psi\rangle^{C}|\psi\rangle^{DEF})=\sqrt{h_{DE}+h_{F}}/2.$
But
$C_{q(2)}(|\psi\rangle^{AB}|\psi\rangle^{C}|\psi\rangle^{DEF})=\sqrt{(h_{A}+h_{B}+h_{D}+h_{E}+h_{F})/6},$
which can not guarantee $C_{q(3)}\leq C_{q(2)}$. In addition,
$\qquad C_{q(3)}(|\psi\rangle^{AB}|\psi\rangle^{C}|\psi\rangle^{DEF})=\sqrt{h_{F}}/2<C_{q(3)}(|\psi\rangle^{AB}|\psi\rangle^{DEF})=\sqrt{2h_{F}/3}$
implies that $C_{q(3)}$ is not coarsening monotonic. For $|\psi\rangle^{AB}|\psi\rangle^{CD}$, $C_{q(2)}(|\psi\rangle^{AB}|\psi\rangle^{CD})=\sqrt{(h_{A}+h_{C})/2}$ may be not larger than $C_{q(2)}(AB|C|D)=\sqrt{2h_{C}/3}$, i.e., it is not tightly coarsening monotonic. Replacing $C_{q(k)}$ with $C_{\alpha(k)}$ in arguments above, we get that $C_{\alpha(k)}$ has the same property as that of $C_{q(k)}$.
Consider $|\psi\rangle^{AB}|\psi\rangle^{CD}$, we have
$C_{G,q(2)}(|\psi\rangle^{AB}|\psi\rangle^{CD})=\sqrt{(h_{A}+h_{C})/2}$
$<$ $C_{G,q(2)}(|\psi\rangle^{AB})+C_{G-q(2)}(|\psi\rangle^{CD})=\sqrt{h_{A}}+\sqrt{h_{C}}$
whenever $h_{A}h_{C}>0$. So it is not additive. Let $|\psi_{1}\rangle=|\psi\rangle^{ABC}|\psi\rangle^{D}$ with $\rho^{A}=\rho^{B}=\rho^{C}$. It turns out that
$C_{G,q(3)}(|\psi_{1}\rangle)=\sqrt[20]{2h_{A}^{10}/9}>C_{G,q(2)}(|\psi_{1}\rangle)=\sqrt{3h_{A}/4}.$
Namely, it is not $k$-monotonic. Let $|\psi_{2}\rangle=|\psi\rangle^{AB}|\psi\rangle^{C}|\psi\rangle^{D}$. Then
$C_{G,q(2)}(|\psi_{2}\rangle)=\sqrt{2h_{A}}/2<C_{G,q(2)}(|\psi\rangle^{AB}|\psi\rangle^{D})$
$=$ $C_{G,q(2)}(A|B|CD)=\sqrt{2h_{A}/3},$
i.e., it is neither coarsening monotonic nor tightly coarsening monotonic. $C_{G,\alpha(k)}$ has the same property as that of $C_{G,q(k)}$.
By definitions, both $E_{(k)}$ and $E^{\prime}_{(k)}$ are not genuine $k$-partite entanglement measures.
5 Examples
[…]
Let $|\psi\rangle=|GHZ_{4}\rangle^{ABCD}|W_{3}\rangle^{EFG}|\psi\rangle^{H}$ with $|GHZ_{4}\rangle^{ABCD}$ is the four-qubit GHZ state and $|W_{3}\rangle^{EFG}$ is the three-qubit W state. Then
$\acute{C}_{(4)}(|\psi\rangle)$ $=$ $2,\quad\acute{C}_{(3)}(|\psi\rangle)=\acute{C}_{(2)}(|\psi\rangle)=2+\sqrt{2},$
$\mathcal{C}_{(4)}(|\psi\rangle)$ $=$ $7/2,\quad\mathcal{C}_{(3)}(|\psi\rangle)=\acute{C}_{(2)}(|\psi\rangle)=7/2+\sqrt{2},$
$C^{\prime}_{(4)}(|\psi\rangle)$ $=$ $3/2,\quad C^{\prime}_{(3)}(|\psi\rangle)=1+2\sqrt{2}/3,\quad C^{\prime}_{(2)}(|\psi\rangle)=2+\sqrt{2},$
$E_{(4)}(|\psi\rangle)$ $=$ $2,\quad E_{(3)}(|\psi\rangle)=E_{(2)}(|\psi\rangle)=1+\frac{3}{2}\log_{2}3,$
$\mathcal{E}_{(4)}(|\psi\rangle)$ $=$ $7/2,\quad\mathcal{E}_{(3)}(|\psi\rangle)=\mathcal{E}_{(2)}(|\psi\rangle)=5/2+\frac{3}{2}\log_{2}3,$
$E^{\prime}_{(4)}(|\psi\rangle)$ $=$ $3/2,\quad E^{\prime}_{(3)}(|\psi\rangle)=1/3+\log_{2}3,\quad E^{\prime}_{(2)}(|\psi\rangle)=1+\frac{3}{2}\log_{2}3.$
For $|\phi\rangle=|W_{3}\rangle|\psi^{+}\rangle=\frac{1}{\sqrt{6}}(|100\rangle+|010\rangle+|001\rangle)(|00\rangle+|11\rangle)$, we have
$\acute{C}_{(3)}(|\psi\rangle)$ $=$ $\sqrt{2},\quad\acute{C}_{(2)}(|\psi\rangle)=1+\sqrt{2},$
$\mathcal{C}_{(3)}(|\psi\rangle)$ $=$ $\sqrt{2},\quad\mathcal{C}_{(2)}(|\psi\rangle)=1+\sqrt{2},$
$C^{\prime}_{(3)}(|\psi\rangle)$ $=$ $2\sqrt{2}/3,\quad C^{\prime}_{(2)}(|\psi\rangle)=1+\sqrt{2},$
$E_{(3)}(|\psi\rangle)$ $=$ $\frac{3}{2}\log_{2}3-1,\quad E_{(2)}(|\psi\rangle)=\frac{3}{2}\log_{2}3,$
$\mathcal{E}_{(3)}(|\psi\rangle)$ $=$ $\frac{3}{2}\log_{2}3-1,\quad\mathcal{E}_{(2)}(|\psi\rangle)=\frac{3}{2}\log_{2}3,$
$E^{\prime}_{(3)}(|\psi\rangle)$ $=$ $\log_{2}3-\frac{2}{3},\quad E^{\prime}_{(2)}(|\psi\rangle)=\frac{3}{2}\log_{2}3.$
6 Conclusion
We have established the axiomatic postulates for complete measure of the $k$-partite entanglement and presented two classes of $k$-partite entanglement measures. Comparing with the axiomatic postulates of complete $k$-entanglement measure in Ref. [2], both the unification condition and the hierarchy condition were modified accordingly in order to make them in consistence with the type of entanglement considered. Together with the complete $k$-entanglement measure, we get a further progress in characterizing of multipartite entanglement. In comparison, although the $k$-PEM is far different from the $k$-entanglement measure, it has some similarities to the $k$-entanglement measure: all of them can be defined by the reduced function and in such a sense the completeness is always related to the subadditivity of the reduced function. In addition, we can discuss the monogamy and the complete monogamy relations of the $k$-entanglement measure, but it seems not compatible for $k$-PEM. Going further, our result is applicable for other $k$-partite measure of quantum correlations since it is based on the coarsening relation of partitions.
This work is supported by the National Natural Science Foundation of China under Grant Nos. 12471434 and 11971277, the Program for Young Talents of Science and Technology in Universities of Inner Mongolia Autonomous Region under Grant No. NJYT25010, and the High-Level Talent Research Start-up Fund of Inner Mongolia University under Grant No. 10000-2311210/049.